How many grams of Strontium (II) nitrate are require to produce .250 liters of a 1.25 M solution?



1
Expert's answer
2021-04-16T05:00:25-0400

Moles = M × V(L)

= 1.25 × 0.250

=0.3125 mol

Mass = moles × MM

= 0.3125 × 211.6

= [66.1] g